At one point, some Treasury bonds were callable. Consider the prices on the following three Treasury issues as of May 15, 2016:

7.05 May 20 n 110.31250 110.37500 .37500 5.39
8.80 May 20 107.43750 107.50000 .12500 5.35
12.55 May 20 145.93750 146.12500 .43750 5.43


The bond in the middle is callable in February 2017. What is the implied value of the call feature? Assume a par value of $1,000. (Hint: Is there a way to combine the two noncallable issues to create an issue that has the same coupon as the callable bond?) (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.)

Calculate Call value $ ________
Answer #1

ANSWER:

As the hint implies, we will first look at the coupon rates. Imagine that we have money to invest, and we want a 8.80% return, but we cannot invest in Bond 2. The other option is to invest part of our money into Bond 1 and part into Bond 3, so that the average rate of return is 8.80%

Formula: Rate 2 = Rate 1 * (X) + Rate 3 * (1-X)

8.80 = 7.05 X + 12.55(1 -X)

8.80 = 7.05 X + 12.55 - 12.55 X

12.55 X - 7.05 X = 12.55 - 8.80

5.50 X = 3.75

X = 3.75 / 5.50

X = 0.681818

Expected Price of Bond 2:

Price 2 = (Price 1 * 0.681818) + (Price 3 * 0.31819)

Price 2 = (110.37500 * 0.681818) + (146.12500 * 0.31819)

Price 2 = (75.25566) + (46.49551)

Price 2 = 121.75117

NOW EXPECTED PRICE - GIVEN PRICE 2:

121.75117 - 107.50000

= 14.25117

ASSUMING BOND VALUES IS $1000

CALL FEATURE IS $ 142.5117